OpenStudy (anonymous):

step by step help 2x+y=-6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

where did 7 come from

OpenStudy (anonymous):

he was wrong what are you trying to find?

OpenStudy (anonymous):

it is system of equations:elimination method

OpenStudy (anonymous):

ok what is the other equation

OpenStudy (anonymous):

3x+y= -10

OpenStudy (anonymous):

2x+y=-6 3x+y=-10 ok, then substract the to eq from each other 2x+y=-6 -(3x+y=-10) so -3x-y=10 so 2x+y=-6 -3x-y=10 -x=4 divide by -1 both side so x=-4 now sub in on of the eq 2(-4)+y=-6 -8+y=-6 +8 to both sides so y=2

OpenStudy (anonymous):

do you follow this?

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

how do you gt the solution suppose to be 2 answers

OpenStudy (anonymous):

it is x=-4 and y= 2

OpenStudy (anonymous):

what site can i use if i need more help?

OpenStudy (anonymous):

this one or ask.com

OpenStudy (anonymous):

or me :-)

OpenStudy (anonymous):

is there a short cut of doing these equations?

OpenStudy (anonymous):

no just look at the two variables and see which is easiest to elimate. some times you have to multiply both equations by numbers to elimate one of the variables, sorry.

OpenStudy (anonymous):

8x-y=20 -5x+y=-8 can you help me with this prob